# When Collaborative Maintenance Falls Short: The Persistence of Retracted Papers on Wikipedia

### v2

@@ 4.2.3 Factors Influencing the Time to Correction: AFT Model Results
The visibility and accessibility of a paper are significantly associated with its correction speed. Papers that are open access are corrected significantly faster than those behind a paywall ($\text{Time Ratio}=0.522$, $p<0.001$). Similarly, a higher level of Twitter/X attention prior to retraction is associated with a faster correction time ($\text{Time Ratio}=0.567$, $p<0.001$), suggesting that public attention on social media may serve as an effective signal to Wikipedia editors.
The academic domain of a paper also shapes the speed of its correction. Citations of papers in the Physical Sciences ($\text{Time Ratio}=0.359$, $p<0.001$) and Social Sciences ($\text{Time Ratio}=0.525$, $p<0.05$) are corrected significantly faster relative to Health Sciences, while citations in the Life Sciences ($\text{Time Ratio}=0.939$, $p=0.769$) show no significant difference with respect to Health Sciences. These findings suggest that domains with greater online or editorial attention may experience more timely corrections, while others may be overlooked. As our descriptive analysis in Figure 4 shows, the Health Sciences domain exhibits the most acute over-representation of retracted work, and we selected it as the reference category for our model. All reported domain effects are therefore relative to this baseline of heightened vulnerability.
Conversely, a higher academic citation count at the start of the risk period (i.e., the point after a paper is both retracted and cited on Wikipedia) is linked to a slower correction ($\text{Time Ratio}=1.352$, $p<0.05$). This may indicate that more highly-cited papers are more deeply embedded in Wikipedia’s content, making editors more hesitant to remove or amend them. The research team size was not found to have a significant effect on the correction time.
@@
Not all retractions are treated equally on Wikipedia. Papers retracted for reasons such as fraud, manipulation, or unreliable results are corrected significantly faster ($\text{Time Ratio}=0.786$, $\textit{p}<0.05$), suggesting that the severity of the retraction notice influences editorial response. When retraction notices explicitly flag misconduct or data concerns, editors may be more likely to notice and address the problematic citation, likely due to heightened visibility or urgency. However, this does not diminish the importance of addressing other types of retractions, such as those resulting from procedural or ethical issues, which may take longer to be recognized or corrected.
Neither the time between a paper’s publication and retraction ($\text{Time Ratio}=0.854$, $p=0.186$) nor the presence of a pre-retraction editorial event ($\text{Time Ratio}=0.899$, $p=0.707$) had a significant effect on the time to correction.
@@
The characteristics of the Wikipedia page hosting the citation are significant predictors of correction speed. A greater number of unique human editors contributing to a Wikipedia page is significantly associated with a faster correction ($\text{Time Ratio}=0.789$, $p<0.01$). This suggests that broader editorial engagement may enhance collective oversight, thereby increasing the likelihood that problematic citations are identified and addressed more promptly.
Furthermore, we find that citations that were flagged as controversial, e.g., those accompanied by dispute tags or appearing in debated content, are corrected significantly faster ($\text{Time Ratio}=0.498$, $p<0.05$). This suggests that early signals of conflict or scrutiny may increase the visibility of problematic content and prompt corrective action. Together, these findings highlight that the scale of human participation and oversight is a key factor associated with the accelerated repair of retracted citations on Wikipedia.
@@ 5 Discussion
The persistence of retracted paper citations on Wikipedia represents a critical vulnerability in the community’s ability to maintain epistemic integrity collaboratively. While Wikipedia is often recognized for its agility in responding to breaking news and high-salience events keegan2012staying ; twyman2017blm , our findings reveal a much slower and uneven response to post-publication knowledge failures. Specifically, 71.6% of the retracted paper citations we examined were problematic. Our survival analysis shows the extent of this delay: the median time to correction is 3.68 years, and a “long tail” of persistence indicates that 67.3% of problematic citations remain uncorrected after nearly 2.74 years. These long delays reflect a gap between Wikipedia’s technical infrastructure for correction and the community’s ability to mobilize editors. Given that Wikipedia continues to serve as a key access point for scientific information, this fragility in information repair raises essential concerns for both end users and the broader knowledge ecosystem.
@@
Our findings can be understood through the lens of a classic coordination challenge in large-scale collaborative systems, where achieving quality is fundamentally dependent on successful coordination between contributors kittur2008harnessing . This is more than a simple information routing issue, where public retraction notices fail to reach the right editors and articles ackerman2013sharing . The challenge is also one of collective sensemaking and action keegan2013hot ; weick2005organizing . Even when editors are aware of a retraction, they may disagree on the appropriate response—whether to remove the citation entirely or add a warning viegas2004studying ; kittur2008harnessing —or on the urgency of the repair compared to other maintenance tasks, as volunteer attention is a scarce resource in large-scale peer production systems kittur2007he . Therefore, the persistence of these citations reveals a breakdown not only in information flow, but also in the community’s ability to achieve consensus and prioritize corrective action. The remainder of this discussion interprets our empirical results through this framework, examining how different factors related to the paper, the retraction event, and the community enhance or inhibit this coordination.
Our analysis reveals that the prevalence of problematic retracted citations is not evenly distributed across academic domains. The Health Sciences domain exhibits the highest rate of retracted paper citations on Wikipedia relative to the total number of retracted papers in that field. In fact, retracted Health Science papers are over three times as likely to be cited on Wikipedia as non-retracted ones. This overrepresentation is particularly troubling given Wikipedia’s prominent role as a public health information source heilman2011wikipedia . While our results show that citations in the Physical and Social Sciences are corrected more quickly, the elevated citation rate in the Health Sciences still increases overall exposure to retracted findings. These findings underscore the necessity for domain-specific strategies to mitigate citation decay and enhance epistemic integrity in high-risk areas, such as health communication.
We also find that not all retractions are treated equally. Citations to papers retracted for high-salience reasons, such as fraud, data manipulation, or serious methodological error, are corrected significantly faster than those withdrawn for ambiguous or technical issues, like authorship disputes or duplication. This suggests that Wikipedia editors may prioritize fixing citations to papers that seem more serious or clearly problematic, such as those involving fraud or significant errors, while not prioritizing less severe or technical retractions. This editing pattern reflects a common trend in collaborative work during crises, where people focus more on problems that seem urgent or clearly defined, and are less likely to act when issues are more complex to recognize or explain twyman2017blm ; keegan2013hot .
A particularly robust finding is the role of human editorial activity. Articles edited by more unique contributors have their retracted citations corrected more quickly. This reinforces established CSCW literature on the benefits of having a larger number of active editors working on an article, which enhances content quality through diversity and redundancy halfaker2013rise ; forte2012coordination . However, it also highlights a structural inequality: many Wikipedia articles receive little editorial attention and thus lack the oversight necessary for timely citation repair. These gaps in attention create a long tail of vulnerability, where outdated or invalid knowledge objects remain visible and unchallenged.
@@
A key consideration is the role of automated tools, such as RetractionBot Headbomb2024Signpost . This bot exemplifies the specialized roles that automated agents play in Wikipedia’s quality control ecosystem zheng2019roles . It primarily serves an editorial audience. By systematically adding a prominent template to the reference section, the bot is highly effective at its specific task of signaling a source’s retracted status to editors engaged in verification and maintenance. The promise of such tools lies in this efficiency for well-defined tasks, and the community has become highly reliant on them for routine maintenance. For instance, the temporary failure of another specialized bot, an anti-vandalism tool, was shown to nearly double the time it took for the Wikipedia community to revert malicious edits geiger2013levee . However, our work highlights a persistent gap between the effectiveness of automation for these specific, often editor-facing tasks and the challenges of repairing more nuanced, epistemic issues for a general reader. This distinction is key: while a bot can efficiently apply a “technical flag,” this action is distinct from the substantive, contextual repair required to update an article’s main text. This reveals a fundamental sociotechnical challenge in designing workflows that successfully translate automated alerts into the complex, collaborative human judgment needed for complete epistemic repair.
Our analysis also points to the role of external visibility cues in prompting correction. Citations to papers that received significant social media attention—particularly on Twitter/X—were corrected more quickly. This finding suggests that social signals external to Wikipedia can help surface retractions to editors and may serve as an informal triage mechanism. Similarly, our finding that open access papers are corrected faster implies that free access can accelerate editorial verification and repair. These findings suggest that social signals and open accessibility can help surface retractions to editors. However, not all forms of visibility lead to faster corrections. We found that papers with a higher academic citation count are actually corrected slower, perhaps because their claims are more accepted and editors are more hesitant to challenge them. These patterns emphasize the importance of salience, not just in the nature of the retraction but in its discoverability.
@@
Together, these findings highlight a multilayered vulnerability in Wikipedia’s citation infrastructure—one rooted not in neglect, but in the subtle interplay between visibility, editorial capacity, and the nature of retraction events. Ultimately, this challenge is not unique to Wikipedia. Similar coordination issues likely exist in other distributed knowledge systems, such as open-source software projects managing bug fixes across dependencies, or collaborative fact-checking platforms striving to correct misinformation that has already propagated widely forte2012coordination ; kittur2008harnessing . Future work should continue to explore how to better surface and prioritize these risks, especially in domains where epistemic failure carries real-world consequences. Our work thus uses Wikipedia as a case study to illuminate a general vulnerability in the maintenance of large-scale, collaboratively produced knowledge, highlighting the critical need for systems that not only detect errors but also facilitate the complex social coordination required to repair them ackerman2013sharing ; jackson201411 .
@@ 5.1 Design Implications
Our findings suggest several design opportunities to address a fundamental coordination challenge of how to efficiently identify and repair citations to retracted scientific papers within Wikipedia’s maintenance ecosystem. The core issue is not a lack of information—retraction notices are public and trackable—but a two-fold coordination challenge: first, a failure in connecting this critical information with the specific contexts and contributors responsible for repair; second, the subsequent challenge of achieving collective sensemaking to agree upon and prioritize a corrective action. Our survival analysis quantifies the consequences of this breakdown, revealing that correction delays can last for years. Crucially, our model demonstrates that signals of human attention may bridge the persistent gap between information availability and coordinated action. The following design directions are therefore aimed at strengthening the sociotechnical mechanisms that facilitate this crucial routing and coordination work.
@@ 5.1.1 Improving Retraction Visibility in Context
Our analysis reveals that early, human-generated warnings are effective; citations flagged as controversial are corrected significantly faster. However, such in-text flags are applied inconsistently. This suggests a need for tools that make these contextual warnings easier to create and more visible. To solve this information routing problem, Wikipedia’s editing interface could surface retraction indicators directly within citation markup or during editing interactions. For example, editors could receive contextual alerts (e.g., a badge or tooltip) when adding or modifying citations to retracted sources, along with links to the official retraction notice. By embedding retraction cues within the editing workflow, these changes would align with CSCW insights about situated action and the importance of interface-level nudges for distributed maintenance work geiger2017operationalizing .
5.1.2 Prioritizing Correction Work through Dashboard Support
Our analysis shows that pages edited by a larger number of unique human contributors have their retracted citations corrected more quickly, reinforcing CSCW research on the benefits of crowd diversity and sustained participation in maintaining information quality forte2012coordination ; halfaker2013rise . However, retracted citations are not evenly distributed: some academic domains—most notably Health Sciences—exhibit disproportionately high rates of retracted paper citations relative to non-retracted ones. While our AFT model shows citations in other domains are corrected faster, their elevated citation rate increases overall exposure to problematic knowledge, particularly in the Health and Life Sciences, which our model found to have similarly slow correction speeds.
These patterns suggest that editorial attention is both influential and unevenly distributed. Wikipedia could address this by integrating retraction data into community coordination tools, such as dashboards or watchlists, to help surface under-attended articles that cite retracted work. To help editors triage their efforts, pages could be prioritized using a risk score model. This model should combine standard metrics, such as article traffic, with the novel predictors identified in our analysis. For instance, our findings suggest that the highest priority should be given to citations of papers with a high academic citation count and paywalled access, especially when they appear in articles with a low number of unique editors, as our AFT model shows these factors are significantly associated with slower correction times. This approach reflects successful CSCW strategies for managing maintenance workloads and prioritizing collaborative repair tasks across distributed systems teblunthuis2018revisiting .
5.1.3 Automating Credibility Cues through Structured Metadata

### v3

@@ 4.2.4 Associations with Time to Correction: Extended Cox Model Results
The visibility and accessibility of a paper are significantly associated with its correction hazard. Consistent with our expectations, papers that are Open Access are corrected at a significantly higher rate than those behind a paywall (HR =1.68=1.68, $p<0.05$). We find no evidence that this association changes over time (interaction $p=0.778$). Similarly, a higher level of Twitter/X attention prior to retraction is linked to a significantly faster correction time (HR =1.62=1.62, $p<0.001$), suggesting that public attention on social media serves as an effective signal to Wikipedia editors. Notably, we found this association to be time-dependent (Interaction HR =0.97=0.97, $p<0.05$): the positive association between tweet volume and correction hazard attenuates as the risk period lengthens.
Conversely, a higher academic citation count at the risk start is significantly associated with a slower correction (HR =0.85=0.85, $p<0.05$). This aligns with the idea that highly cited papers may possess an “authority shield” that creates editorial inertia, making them harder to correct even after retraction.
@@
Time to retraction (publication-to-retraction duration) does not show a statistically significant association with correction timing ($p=0.063$). In addition, the presence of a pre-retraction editorial event (e.g., an expression of concern) is not significantly associated with correction timing ($p=0.921$). Retraction reason does not show a constant (time-invariant) association ($p=0.881$); however, its association varies over time (interaction HR =1.07=1.07, $p<0.05$), indicating that differences by reason become more pronounced later in the risk period. Overall, these results suggest that retraction-event characteristics show limited average associations with reader-facing correction timing, with some effects emerging in a time-dependent manner.
@@
Not all retractions are treated equally on Wikipedia. Citations flagged by RetractionBot exhibited the highest hazard of correction (HR =2.34=2.34, $p<0.001$), highlighting the role of bot-assisted maintenance in accelerating reader-facing repair.
We find that the structural organization of a Wikipedia page is a more reliable predictor of the correction hazard than the sheer volume of editorial activity. We used categories count as a proxy for a page’s integration into Wikipedia’s knowledge graph and found that retracted citations on pages with more categories are associated with a significantly higher hazard of correction (HR =1.27=1.27, $p<0.001$).
Explicit signals of dispute were also strongly associated with the hazard of correction. Citations that were flagged as controversial at the time of addition had a significantly higher hazard of correction (HR =1.63=1.63, $p<0.05$). This suggests that visual cues of disagreement successfully attract attention to specific bibliographic errors. However, none of the number of unique editors ($p=0.436$), the number of media files ($p=0.614$), or the number of references ($p=0.534$) was significant. This indicates that simply having a larger crowd of general contributors does not guarantee the specific form of scrutiny required to identify and fix retracted citations.
@@ 5 Discussion
The persistence of retracted paper citations on Wikipedia highlights a coordination vulnerability in how epistemic risks are handled within a large-scale coordination platform. Rather than directly observing editorial deliberation or decision-making, our analysis examines the outcomes of collaborative maintenance as recorded through revision histories. From this system-level perspective, retracted citations serve as an indicator of how effectively signals of epistemic failure are surfaced, routed, and translated into reader-facing repair across human editors, automated tools, and platform infrastructure. While Wikipedia is often recognized for its agility in responding to breaking news and high-salience events Keegan et al. (2012); Twyman et al. (2017), our findings reveal a much slower and uneven response to post-publication knowledge failures. Specifically, 71.6% (845/1,181) of the retracted paper–page citation pairs were initially problematic—i.e., they were introduced before retraction or added after retraction without an in-text acknowledgment. By the end of the observation period, 32.9% of these problematic citations remained uncorrected (278/845), while the rest were corrected through removal or in-text acknowledgment. Our survival shows that the median time to correction is 3.68 years, and a “long tail” of persistence indicates that (by the Kaplan–Meier estimate) 67.3% of problematic citations remain uncorrected after approximately 1,000 days (2.74 years). These long delays reflect a gap between Wikipedia’s technical infrastructure for correction and the community’s ability to mobilize editors. Given that Wikipedia continues to serve as a key access point for scientific information, this fragility in information repair raises essential concerns for both end users and the broader knowledge ecosystem.
@@
Our findings can be understood through the lens of a classic coordination challenge in large-scale collaborative systems, where achieving quality is fundamentally dependent on successful coordination among contributors Kittur and Kraut (2008). This is more than a simple information routing issue, where public retraction notices fail to reach the right editors and articles Ackerman et al. (2013). The challenge is also one of collective sensemaking and action Keegan et al. (2013); Weick et al. (2005). Even when editors are aware of a retraction, they may disagree on the appropriate response—whether to remove the citation entirely or add a warning Viégas et al. (2004); Kittur and Kraut (2008)—or on the urgency of the repair compared to other maintenance tasks, as volunteer attention is a scarce resource in large-scale peer production systems Kittur et al. (2007). Accordingly, the long-tail persistence we observe is consistent with coordination frictions at multiple stages of repair: making the issue salient, agreeing on a reader-facing remedy, and allocating scarce attention to execute it. At the same time, our results suggest that these frictions are not uniform: when a retracted citation becomes more visible and actionable (e.g., through targeted signaling), correction occurs substantially sooner, indicating that coordination can succeed once the problem is concretely surfaced and routed. Therefore, the persistence of these citations reveals a breakdown not only in information flow but also variability in the community’s ability to converge on a corrective strategy and prioritize action across pages and time. The remainder of this discussion interprets our empirical results through this framework, examining how different factors related to the paper, the retraction event, and the community are associated with this coordination.
Our analysis reveals that the prevalence of problematic retracted citations is not evenly distributed across academic domains. The Health Sciences domain exhibits the highest rate of retracted paper citations on Wikipedia relative to the total number of retracted papers in that field. In fact, retracted Health Science papers are over three times as likely to be cited on Wikipedia as non-retracted ones. This overrepresentation is particularly troubling given Wikipedia’s prominent role as a public health information source Heilman et al. (2011). Beyond prevalence, our Kaplan–Meier curves stratified by domain suggest that these domains also differ significantly in their correction time (see Figure 7). Citations in the Physical Sciences are corrected most rapidly, with a median persistence of approximately 228 days. In contrast, citations in the Life Sciences and Health Sciences display significantly longer persistence, with median correction times of 2,720 days and 1,846 days, respectively. The Social Sciences show intermediate behavior with a median of 724 days, though estimates in this domain should be interpreted cautiously given the smaller sample size. These results complicate a simple account of domain-specific collaborative maintenance. Specialized communities of practice, such as WikiProject Medicine Heilman et al. (2011), play an essential role in coordinating high-quality biomedical content. Yet, the Health Sciences still exhibit relatively slow correction rates. One plausible interpretation is that this accumulation reflects maintenance debt rather than systematic differences in source selection. Because retractions often occur long after papers are cited on Wikipedia, the ongoing task of monitoring previously accepted sources may exceed the capacity of manual editorial workflows alone. From this perspective, slower correction times in a domain with robust governance structures highlight the limits of relying on purely manual monitoring at scale and motivate the technical interventions discussed in Section 5.1.
We also find that not all retractions are treated equally. In the extended Cox model, we do not observe a stable average difference in correction timing by salient retraction reasons at baseline. However, this association is time-dependent: the reason effect increases over the risk period, indicating that differences by retraction reason become more pronounced later rather than immediately after a citation enters the risk set. This suggests that citations to papers retracted for more salient reasons may be increasingly likely to be corrected as time passes, consistent with a delayed sensemaking process in which “clearer” epistemic failures become easier to justify and act upon after the retraction has had time to circulate. This editing pattern reflects a common trend in collaborative work during crises, in which people focus more on problems that seem urgent or clearly defined and are less likely to act when issues are more complex to recognize or explain Twyman et al. (2017); Keegan et al. (2013).
A critical and counterintuitive finding is the limit of the “wisdom of the crowd” in this specific context. Contrary to the expectation that “given enough eyeballs, all bugs are shallow,” we found that the number of unique editors was not a significant predictor of correction speed. Simply having a larger volume of contributors does not accelerate the repair of deep bibliographic errors. Instead, we found that structural integration matters: citations on pages with a higher categories count are corrected significantly faster. This suggests that for maintenance tasks, organizational structure triumphs over raw participation volume Kittur and Kraut (2008); Halfaker et al. (2013). Pages that are well-categorized are likely better integrated into Wikipedia’s ontology, making them more visible to experienced editors and WikiProjects who perform systematic cleanup Forte et al. (2012); Kittur and Kraut (2008). Furthermore, explicit signals of dispute proved highly effective. Citations flagged as controversial at the time of addition had a significantly higher hazard of correction. Unlike general editor volume, visual cues of disagreement successfully attract attention to specific bibliographic errors, surfacing them from the background noise of page history.
@@
A key consideration is the role of automated tools, such as RetractionBot Headbomb (2024). This bot exemplifies the specialized roles that automated agents play in Wikipedia’s quality control ecosystem Zheng et al. (2019). It primarily serves an editorial audience. By systematically adding a prominent template to the reference section, the bot is highly effective at its specific task of signaling a source’s retracted status to editors engaged in verification and maintenance. The promise of such tools lies in this efficiency for well-defined tasks, and the community has become highly reliant on them for routine maintenance. For instance, the temporary failure of another specialized bot, an anti-vandalism tool, was shown to nearly double the time it took for the Wikipedia community to revert malicious edits Geiger and Halfaker (2013). Similarly, Nyayachavadi et al. Nyayachavadi et al. (2022) demonstrated the community’s dependence on InternetArchiveBot to manage technical decay, such as ‘permanently dead’ links. Our results powerfully validate this reliance in the context of epistemic repair. We found that citations flagged by RetractionBot had the highest hazard of correction, outperforming all other factors. This finding helps resolve the sociotechnical challenge of translating automated alerts into human judgment: it suggests that the “technical flag” applied by the bot serves as an effective scaffold for the substantive, contextual repair required to update the main text. Taken together, our findings suggest a two-stage process: automated monitoring helps surface candidates for review, and human/editorial work performs the reader-facing repair. The bottleneck may lie in coverage-which retracted citations are surfaced for attention—rather than in editors’ willingness to act once a credible signal is present.
Our most counterintuitive finding further underscores the necessity of better workflow design: 20.1% of problematic citations were introduced after the paper had already been officially retracted. This phenomenon persists despite clear publisher-level signals, such as prominent “RETRACTED” watermarks on journal websites. We speculated that this disconnect stems from specific gaps in editorial workflows that allow editors to inadvertently bypass these warnings. For instance, editors may rely on locally stored PDFs downloaded prior to retraction, or engage in secondary citation by copying bibliographic details from other academic works without verifying the sources online. In these scenarios, the editor does not visit the publisher’s landing page. While the failure to repair older citations might be attributed to passive under-maintenance, the active introduction of invalidated knowledge highlights a critical vulnerability: relying solely on external publisher signals is insufficient. This necessitates platform-level interventions detailed in Section 5.1, such as embedding contextual alerts within the editing interface to detect these errors at the time of citation.
Our analysis also points to the role of external visibility cues in prompting correction. Citations to papers that received significant social media attention, particularly on Twitter/X, were associated with a significantly higher hazard of correction. This reflects the broader influence of public visibility on editorial prioritization: while social media activity can serve as a direct signal surfacing specific retractions Zheng et al. (2025), it also acts as a proxy for a paper’s general prominence. High-profile papers naturally attract greater scrutiny across the web, which may increase the probability that their retraction is noticed and addressed on Wikipedia. Crucially, our modeling reveals that this “visibility advantage” is ephemeral. We found a significant negative interaction with time, indicating that the strong initial association between tweet volume and correction hazard diminishes rapidly. In contrast, Open Access availability appears to provide a more sustained structural advantage. This aligns with recent scholarship by Yang et al. Yang et al. (2024), who found that open access reduces barriers to access for volunteer editors. Furthermore, studies by Zheng et al. Zheng et al. (2024); Zheng and Fu (2024) demonstrate that Gold Open Access accelerates the “academic purification” process within the scientific community itself, suggesting that accessibility can facilitate earlier detection and “purification” upstream. Taken together, these evidences are consistent with a broader “transparency advantage”: accessibility may lower informational friction both in formal correction processes and in downstream venues like Wikipedia, although our observational design does not isolate a specific causal pathway. Conversely, not all forms of visibility facilitate repair: papers with higher academic citation counts associated with a lower hazard of correction.This is perhaps because their claims are more accepted, and editors are more hesitant to challenge them. These patterns emphasize the importance of salience: it is not just the existence of a retraction that matters, but its discoverability and the willingness of the community to act upon it.
@@
Together, these findings highlight a multilayered vulnerability in Wikipedia’s citation infrastructure—not a simple story of individual neglect, but a coordination problem shaped by how epistemic risk signals become (or fail to become) discoverable and actionable. This vulnerability emerges from the interplay between visibility cues, structural organization (rather than sheer editor volume), bot-mediated flagging and maintenance workflows, and the salience of retraction events. Ultimately, this challenge is not unique to Wikipedia. Similar coordination issues likely exist in other distributed knowledge systems, such as open-source software projects managing bug fixes across dependencies, or collaborative fact-checking platforms striving to correct misinformation that has already propagated widely Forte et al. (2012); Kittur and Kraut (2008). Future work should continue to explore how to better surface and prioritize these risks, especially in domains where epistemic failure carries real-world consequences. Our work thus uses Wikipedia as a case study to illuminate a general vulnerability in the maintenance of large-scale, collaboratively produced knowledge, highlighting the critical need for systems that not only detect errors but also facilitate the complex social coordination required to repair them Ackerman et al. (2013); Jackson (2014).
@@ 5.1 Design Implications
Our findings suggest several design opportunities to address a fundamental coordination challenge: efficiently identifying and repairing citations to retracted scientific papers within Wikipedia’s maintenance ecosystem. The core issue is not a lack of information—retraction notices are public and trackable—but a two-fold coordination challenge: first, a failure in connecting this critical information with the specific contexts and contributors responsible for repair; second, the subsequent challenge of achieving collective sensemaking to agree upon and prioritize a corrective action. Our survival analysis quantifies the consequences of this breakdown, revealing that correction delays can last for years. Crucially, our model demonstrates that signals of human attention may bridge the persistent gap between information availability and coordinated action. The following design directions are therefore aimed at strengthening the sociotechnical mechanisms that facilitate this crucial routing and coordination work.
@@ 5.1.1 Improving Retraction Visibility in Context
Our analysis reveals that early, human-generated warnings are effective; citations flagged as controversial are corrected significantly faster. However, such in-text flags are applied inconsistently. This suggests a need for tools that make these contextual warnings easier to create and more visible. To solve this information routing problem, Wikipedia’s editing interface could surface retraction indicators directly within citation markup or during editing interactions.
Ideally, this functionality could be integrated into existing MediaWiki services. For instance, the “Citoid” service, which automatically retrieves citation metadata from URLs and DOIs, could be improved to check the retraction status of sources during metadata retrieval automatically. Furthermore, by leveraging the recent deployment of “Edit Check,” which prompts users to add citations for new claims, a similar “Retraction Check” mechanism could be implemented to verify the status of added sources in real time. By embedding retraction cues within the editing workflow, these changes would align with CSCW insights about situated action and the importance of interface-level nudges for distributed maintenance work Geiger and Halfaker (2017).
5.1.2 Prioritizing Structure over Traffic in Maintenance Tools
Our analysis challenges the assumption that high-traffic or highly edited pages are inherently self-correcting. We found that the number of unique editors was not a significant predictor of repair speed, suggesting that “many eyes” do not necessarily spot deep bibliographic errors. Instead, the significance of Categories Count indicates that structural integration is strongly asociated with maintenance. Pages that are well-categorized are likely more visible to WikiProjects and experienced “gnome” editors who use category-based tools to perform systematic cleanup.
This has direct implications for the design of community dashboards and task lists. Current tools often prioritize pages based on traffic or recent activity. Our findings suggest a different approach: maintenance tools should prioritize structural isolation. Algorithms could flag retracted citations on pages with low category counts (i.e., “orphaned” from the maintenance ontology) and high-risk features (e.g., Health Sciences domain, high citation count). By directing attention to these structurally disconnected pages—which our model shows are prone to neglect—dashboards can help extend coverage and route attention to structurally disconnected pages, where correction is otherwise likely to lag TeBlunthuis et al. (2018).
5.1.3 Automating In-Text Credibility Cues through Structured Metadata